# of the a an and or to from in as about but than then so through

### v1

@@ Abstract.
To close that knowledge gap, this paper presents the first large-scale empirical study of floating-point arithmetic usage in statically typed languages across public GitHub repositories. We follow state-of the art mining practices including random sampling and filtering based on only intrinsic properties to avoid bias, and identify floating-point usage by searching for keywords in the source code, and programming language constructs (e.g., loops) by parsing the code. Our evaluation supports the claim often made in papers that floating-point arithmetic is widely used. Comparing statistics such as size and usage of certain constructs and functions, we find that benchmarks used in literature to evaluate automated reasoning techniques for floating-point arithmetic are in certain aspects representative of ‘real-world’ code, but not in all.
We aim for our study and dataset to help future techniques for floating-point arithmetic to be designed and evaluated to match actual users’ expectations.
copyright: none
@@ 1. Introduction
Numerical software is widely used across different domains such as embedded systems, scientific computing, and machine learning. Numerical software frequently makes use of floating-point arithmetic to efficiently approximate computations with real numbers. Some approximation of the infinitely precise reals is necessary, which introduces rounding errors at most arithmetic operations as well as potentially the special values infinity and Not-a-Number (NaN). Overall this make reasoning about the correctness of such software unintuitive, error prone and time consuming. — This could be the beginning of a paper on automated reasoning for floating-point programs, e.g., with a static (Darulova et al., 2018; Solovyev et al., 2018) or a dynamic approach (Zou et al., 2019; Guo and Rubio-González, 2020; Zou et al., 2022a). Such an approach would typically be evaluated either using a small benchmark suite (Damouche et al., 2017), or on a relatively small set of benchmarks hand-picked specifically for that paper.
@@
While hand-picking benchmarks can be helpful to demonstrate particular features and issues, it is unclear how they generalize to “real-world” code. To the best of our knowledge, no large-scale code study has investigated how representative the currently used benchmarks are of “real-world” usage, e.g., in terms of size, operations performed, or frequency of library function usage. Neither do we know how widely used floating-point arithmetic really is.
@@
While real-world floating-point usage is largely anecdotal, the difficulty in developing automated reasoning techniques for floating-point arithmetic is well documented. For example, static analysis techniques bounding worst-case rounding errors (Titolo et al., 2025; Darulova et al., 2018; Solovyev et al., 2018) are limited to small programs (mostly individual functions) and have no or limited support for common program constructs such as conditionals and loops. This is also reflected in the standard benchmark set used in this domain (Damouche et al., 2017). Dynamic approaches aiming to identify specific inputs that induce large rounding errors (Guo and Rubio-González, 2020; Zou et al., 2019), or for repairing such errors (Zou et al., 2022b) are evaluated on different, hand-picked sets of benchmarks, most of which are typically small (e.g., individual functions). In summary, the current benchmarks used align with the current capabilities of the tools, but it is unclear to what extent they align with actual users’ expectations.
@@
This paper presents the first large-scale study of statically typed real-world floating-point code. The aim is to inform current and future research efforts in this domain, for example by aiding (more) representative selections of benchmarks for evaluation, and to guide the development of future benchmarks and reasoning techniques in a practically relevant direction.
@@
Characterising “real-world floating-point code” is non-trivial for several reasons. First, limiting ourselves to publicly available code, real-world code bases use a multitude of programming languages and the volume is too large to be processed and inspected in full. The latter can be addressed by sampling, but to avoid measurement bias, the code must be checked for code clones (Lopes et al., 2017), and filtered to exclude projects that do not contain any code at all and “toy” projects (e.g., from class assignments, or projects that were never updated after the initial commit). While a convenient and continuously updated dataset of GitHub projects aimed for code studies (Dabic et al., 2021) exists, it includes only projects with at least ten stars (to reduce the amount of data). Such filtering has been shown to skew the results significantly (Maj et al., 2024), making it unsuitable to generalise from.
Second, once we have a set of “interesting” projects, there is no existing tool that can efficiently identify and analyze floating-point code. Such analysis must be automated to be able to handle sufficiently large volumes of code. Furthermore, it should be able to handle different programming languages.
@@
[…]
We design and implement a large-scale mining methodology (Section˜3) for detecting and analyzing floating-point usage in statically typed languages at the project, file and function level. Our approach combines intrinsic property filtering, keyword-based detection, fuzzy duplicate elimination and parsing to produce a reliable dataset. Large parts of the methodology and implementation are not specific to floating-point analysis and we expect them to be reused in other large-scale empirical code studies.
[…]
We analyse 453 891453\,891 GitHub projects (Section˜4) and present the results in Section˜5. Our analysis shows that, with 95 % confidence, over 62 % of the projects contain floating-point code, supporting the claim that floating-points are widely used. We are the first to draw this conclusion based on a large-scale code study of open-source code. Most functions in our final dataset tend to be small, and function calls and conditional statements appear more often than loops and (explicitly mentioned) special values. The FPBench (Damouche et al., 2017) benchmarks used in literature to evaluate floating-point techniques show different characteristics, e.g., involving fewer conditionals and more transcendental library function calls. Functions from the GNU Scientific Library, often used in evaluations, appear rarely in our dataset.
[…]
We release a dataset of 10 million real-world floating-point functions extracted from our study  [^1] Available at https://doi.org/10.5281/zenodo.17055623. This dataset sets a new bar in terms of representativeness and scale, and can serve as a foundation for constructing new benchmarks and evaluating floating-point analysis tools.
@@ 2. Related Work
Most works that focus on numerical programs are evaluated using different sets of hand-picked programs. Hand-picking programs can be good for stressing certain types of problems, but can just as easily—intentionally or not—suppress other problems and challenges. It is also unclear how to generalise from hand-picked problems to code “in the real world”.
@@
One benchmark set specifically for floating-point analysis exists (FPBench (Damouche et al., 2017)), and it is typically used to evaluate static analysis tools. However, these benchmarks were hand-picked for individual papers and are typically just arithmetic expressions, and therefore do not likely represent usage or challenges found in real-world programs. Techniques that target programs outside of this set choose their own set of benchmarks, e.g., programs that are larger (Das et al., 2020) or that use specific data structures (Isychev and Darulova, 2023).
@@
Some techniques are evaluated on (a small number of hand-picked functions from) numerical libraries, e.g., the GNU Scientific Library (GSL) (Miao et al., 2024; Guo and Rubio-González, 2020; Zou et al., 2019, 2022b), occasionally extended by some additional functions. Di Franco et al. (Di Franco et al., 2017) study bug characteristics in numerical libraries (NumPy, LAPACK, GSL, etc.). Liew et al. (Liew et al., 2017) have two teams independently hand-pick benchmarks for evaluating symbolic execution. In the HPC community, there are examples of picking a specific proxy application, such as Lulesh, as a representative of “HPC code” (Wang and Rubio-González, 2024). In what way these selected benchmarks are representative of real-world code, e.g., beyond just library code, is unclear and not discussed in these papers.
[…]
@@
We are not aware of any large-scale code study that investigates characteristics of real-world floating-point code. Previous works have investigated how different features are used in practice in different programming languages, for example in the context of static analysis for the R programming language (Sihler et al., 2024), dynamic features in Smalltalk (Callaú et al., 2011), the use of eval and dynamic behavior in JavaScript (Richards et al., 2011, 2010), dynamic features or code complexity in Python (Åkerblom et al., 2014; Grotov et al., 2022), inheritance, method chaining, and streams in Java (Tempero et al., 2008; Dyer et al., 2014; Khatchadourian et al., 2020; Keshk and Dyer, 2023), or gradual types in TypeScript and Python (Troppmann et al., 2024). While this paper also studies features of code, these are sufficiently different that we cannot straight-forwardly re-use existing infrastructure.
While some features in dynamically typed programming languages such as Python can be detected statically (Peng et al., 2021), others (for instance types) require programs to be run on suitable inputs (Åkerblom et al., 2014), which cannot be straight-forwardly automated. Detecting floating-point operations in a language like Python would equally require executing the programs due to typical overloading of arithmetic operators such as addition (+). Since this is not possible to automate at a larger scale, we focus on statically typed programming languages only.
[…]
@@
Our methodology draws inspiration from state-of-the-art practices in GitHub mining. Prior studies show that the majority of GitHub repositories are inactive, personal, or not software projects at all, and that most of the files are duplicates of others (Kalliamvakou et al., 2014; Lopes et al., 2017). Other work demonstrates that filtering repositories using extrinsic characteristics such as the number of stars introduces systematic biases in software repository studies (Borges and Tulio Valente, 2018; Maj et al., 2024). In a reproduction study, Berger et al. reveal the the risks of careless keyword searches and emphasize the necessity of manual validation to ensure methodological soundness (Berger et al., 2019).
@@
Several frameworks exist to simplify code studies on GitHub, but none fully satisfy our requirements. Queryable datasets mined from GitHub exist, yet they are either outdated (Dyer et al., 2015), affected by inconsistencies (Gousios and Spinellis, 2012), or rely on GitHub stars (Dabic et al., 2021; Dabić et al., 2024). World of Code (Ma et al., 2021) continuously aggregates repositories from multiple sources and could be used to implement parts of our methodology, but it does not expose part of the metadata we require and does not allow downloading of full projects. Other frameworks target different aspects of repository analysis. CodeDJ (Maj et al., 2021) ensures reproducible queries over projects by downloading full project histories, which exceeds our needs and would lead to an excessive amount of data being downloaded; the analysis would also still need to be implemented "from scratch". GitCProc (Casalnuovo et al., 2017) enables keyword searches in commit histories but supports only a limited set of languages and does not exclude comments or string literals, which compromise validity. CodeQL (git, 2025) offers semantic analysis across multiple languages but requires full-project access, raising scalability concerns. Finally, although our methodology and implementation use GitHub as the primary data source, they are not specific to it. Both can be extended or adapted to other hosting platforms.
@@ 3. Methodology
Our goal is to study the prevalence and type of code operating on floating-point numbers in “real-world” code. To do so, we design a methodology that mainly follows the approach suggested by Maj et al. (Maj et al., 2024) and that is designed in accordance with the ACM SIGSOFT Empirical Standards for repository mining (Ralph et al., 2021). We first give an overview here, before discussing details in the next subsections.
3.1. Overview
We first define our target population: functions from non-trivial projects hosted on GitHub containing floating-point computations within code written in statically typed languages. The restriction to statically typed code on GitHub is due to practical constraints. GitHub alone contains an enormous amount of public repositories, and we assume here that they form a representative sample of real world software. Focusing on statically typed languages allows us to automate the analysis of code and thus to scale to significantly more projects. Our final dataset consists of individual functions as most floating-point reasoning tools work on a per-function basis, though the corresponding files and projects are easily identified. Our quantitative analysis considers per-project, per-file and per-function properties.
We cannot query GitHub for non-trivial projects containing such functions directly in part because each GitHub API call provides only limited information. We thus design a multi-step selection process or pipeline that progressively reduces the amount of data, increasing the level of granularity from the repository to the file and ultimately to the function level. An overview of the data selection pipeline is shown in Fig.˜1.
We begin by uniformly sampling public repositories on GitHub using GitHub’s Rest API (step 1), and by excluding forks and ‘toy’ projects (step 2 and 3) (defined in Section˜3.2). For each remaining repository, we query the list of languages and retrieve the hash of the last commit on the main branch (step 4). We discard projects that do not have any files written in statically typed languages (step 5). At the end of this step, we have metadata about the repositories, after preliminary filtering.
Next, we download the content of the selected repositories and inspect individual files within these that are written in statically typed programming languages (step 6). For these we identify floating-point usage using “keywords” in the source code directly, e.g., focusing on types (e.g., double and real) or mathematical functions (e.g., sin and exp). We discard data that we do not consider: files that are not written in statically typed languages, files that do not contain any keyword related to numerical computations, and files that are exact duplicates of other files (in any considered repository, step 7). We then parse the remaining files, remove comments, and extract individual functions (step 8). We retain only functions that include keywords related to numerical computations (step 9). In the final step we quantify usage of programming language constructs such as loops or conditionals statements fully automatically based on the parse trees.
The pipeline is deliberately designed to be modular and extensible. The sampling and filtering steps earlier in the pipeline are not specific to floating-point arithmetic and can be re-used for other code studies. We envision the floating-point specific dataset collected at the end of the pipeline to be used, e.g., to sample or validate benchmarks with specific features that are used for evaluating new floating-point automated reasoning techniques.

### v2

# arXiv:2509.04936v2[cs.PL] 27 Feb 2026

@@ Abstract.
To close that knowledge gap, this paper presents the first large-scale empirical study of floating-point arithmetic usage across public GitHub repositories. We focus on statically typed languages to allow our study to scale to millions of repositories. We follow state-of the art mining practices including random sampling and filtering based on only intrinsic properties to avoid bias, and identify floating-point usage by searching for keywords in the source code, and programming language constructs (e.g., loops) by parsing the code. Our evaluation supports the claim often made in papers that floating-point arithmetic is widely used. Comparing statistics such as size and usage of certain constructs and functions, we find that benchmarks used in literature to evaluate automated reasoning techniques for floating-point arithmetic are in certain aspects representative of ‘real-world’ code, but not in all.
We publish a dataset of 10 million real-world floating-point functions extracted from our study. We demonstrate in a case study how it may be used to identify new floating-point benchmarks and help future techniques for floating-point arithmetic to be designed and evaluated to match actual users’ expectations.
floating-point arithmetic, large-scale code analysis, repository mining
$c$s: Software and its engineering Software libraries and repositoriesccs: Software and its engineering Automated static analysisccs: Information systems Data miningcopyright: ccdoi: 10.1145/3798203journalyear: 2026journal: PACMPLjournalvolume: 10journalnumber: OOPSLA1article: 95publicationmonth: 4
@@ 1. Introduction
Numerical software is widely used across different domains such as embedded systems, scientific computing, and machine learning. Such software frequently makes use of floating-point arithmetic to efficiently approximate computations with real numbers. Some approximation of the infinitely precise reals is fundamentally and practically necessary, and introduces rounding errors at most arithmetic operations as well as potentially the special values infinity and Not-a-Number (NaN). Overall this makes reasoning about the correctness of such software unintuitive, error prone and time consuming.
This could be the beginning of a paper on automated reasoning for floating-point programs, e.g., with a static (Darulova et al., 2018; Solovyev et al., 2018) or a dynamic approach (Zou et al., 2019; Guo and Rubio-González, 2020; Zou et al., 2022). Such an approach would typically be evaluated either using a small benchmark suite (Damouche et al., 2017), or on a relatively small set of benchmarks hand-picked specifically for that paper.
@@
While hand-picked benchmarks can be helpful to demonstrate particular features and issues, it is unclear how they generalise to “real-world” code. To the best of our knowledge, no large-scale code study has investigated how representative the currently used floating-point benchmarks are of “real-world” usage, e.g., in terms of size, operations performed, or frequency of library function usage. Neither do we know how prevalent the use of floating-point arithmetic is in modern software.
@@
While real-world floating-point usage is anecdotal, the difficulty in developing automated reasoning techniques for floating-point arithmetic is well documented. For example, static analysis techniques bounding worst-case rounding errors (Titolo et al., 2025; Darulova et al., 2018; Solovyev et al., 2018) are limited to small programs (mostly individual functions) and have no or limited support for common program constructs such as conditionals and loops. This is also reflected in the standard benchmark set (Damouche et al., 2017) used in this domain. Dynamic approaches aiming to identify specific inputs that induce large rounding errors (Guo and Rubio-González, 2020; Zou et al., 2019), or for repairing such errors (Zou et al., 2022; Panchekha et al., 2015) are evaluated on different, hand-picked sets of benchmarks, most of which are typically small (e.g., individual functions). In summary, the current benchmarks used align with the current capabilities of the tools, but it is unclear to what extent they align with actual users’ expectations in terms of the code they want the tools to reason about. While community benchmarks are a great way of focusing research, unless the benchmarks are “relevant”, they may end up directing research in the wrong direction.
@@
This paper presents the first large-scale study of real-world floating-point code. The aim is to inform current and future research efforts in this domain, for example by aiding (more) representative selections of benchmarks for evaluation, and to guide the development of future benchmarks and reasoning techniques in a practically relevant direction. With that in mind, we address the following research questions:
  • RQ1. How prevalent is floating-point arithmetic in open-source statically typed code on GitHub?
  • RQ2. What are the characteristics of floating-point code in terms of programming language constructs and size?
  • RQ3. How representative are benchmarks used in the literature to evaluate floating-point analyses of real-world floating-point code?
@@
Characterising “real-world floating-point code” is non-trivial for several reasons. First, limiting ourselves to publicly available code, real-world code bases use a multitude of programming languages and the volume is too large to be processed and inspected in full. The latter can be addressed by sampling, but to avoid measurement bias, the code must be checked for code clones (lopesDejaVuMapCode2017), and filtered to exclude projects that do not contain any code at all and “toy” projects (e.g., projects that were never updated after the initial commit). While a convenient and continuously updated dataset of GitHub projects (dabicSamplingProjectsGitHub2021) aimed for code studies as ours exists, it includes only projects with at least ten stars (to reduce the amount of data). Such filtering has been shown to skew the results significantly (majFaultOurStars2024), making it unsuitable to generalise from.
Second, once we have a set of “interesting” projects, there is no existing tool that can efficiently identify and analyse floating-point code. To obtain representative results, such an analysis must support a multitude of programming languages and be automated to be able to handle sufficiently large volumes of code. Our insight is that in statically typed languages the type annotations on e.g., function parameters allow us to identify floating-point code with an inexpensive and largely programming language-agnostic approach by “grepping” for suitable keywords in source code. To identify floating-point code, type annotations are critical; it is not enough to look for the presence of arithmetic operators (e.g., +) as these are often overloaded to work for integers, strings or other types. This is unlike other code properties such as imports of specific libraries that can be unambiguously identified even in dynamically typed source code. [^1] Our study shows that mathematical library functions (e.g., sin, exp) cannot be used to identify floating-point code either, as much floating-point code does not use them. In this paper, we therefore limit ourselves to studying floating-point usage in statically typed languages.
Contributions
@@
Large-Scale Floating-Point Code Mining
We design and implement a large-scale mining methodology (Section˜3) for detecting and analysing floating-point usage in statically typed languages at the project, file and function level. Our approach is fully automated and combines random sampling, project filtering based on intrinsic properties, and fuzzy duplicate elimination to produce a reliable dataset.
Our methodology identifies floating-point usage in statically typed programming languages using keywords in the source code directly. This allows us to consider all statically typed languages that are recognised by GitHub in a scalable and fully automated way, as long as these languages have floating-point types and a language reference. For the most commonly used languages (covering more than 92 % of the sampled code), we parse the code to identify programming language constructs (e.g., loops) using a language-agnostic parsing library.
Large parts of the methodology and implementation are not specific to floating-point analysis (nor to statically typed code, although we rely on types for identifying relevant code) and we expect them to be reusable in other large-scale empirical code studies. The code is available as open-source [^2] Available at: https://github.com/fxpl/scyros..
Quantitative Analysis
We analyse 447 209447\,209 (randomly sampled) GitHub projects (Section˜4) and present the results in Section˜5. Our analysis shows that, with 95 % confidence, over 62 % of the projects contain floating-point code, supporting the often made claim that floating-point numbers are widely used. While this confirms folklore beliefs, we are the first to draw this conclusion based on a large-scale code study of open-source code.
Most functions in our final dataset tend to be small, and function calls and conditional statements appear more often than loops and (explicitly mentioned) special values. The FPBench (Damouche et al., 2017) benchmark suite used in literature to evaluate floating-point techniques shows different characteristics, e.g., involving fewer conditionals and more transcendental library function calls. Functions from the GNU Scientific Library, also sometimes used in evaluations, appear rarely in our dataset.
To the best of our knowledge, our study is the first to provide a clear picture of how floating-point numbers are actually used in real-world code, and indicates that currently used floating-point benchmarks are largely not representative of such code.
Floating-Point Functions Dataset and Challenge Benchmarks
For the benefit of other researchers, we release a dataset of 10 million real-world floating-point functions extracted from our study [^3] Available at https://doi.org/10.5281/zenodo.17055622.. This dataset focuses on functions, since most static and dynamic floating-point reasoning tools today work on a per-function basis, so this dataset is most (immediately) relevant for their benchmarking. We intend this dataset to serve as a foundation for constructing new realistic benchmarks and evaluating floating-point reasoning tools in a practically relevant way, and invite others to contribute in this undertaking.
We show how our corpus can be used to generate realistic, ready-to-run benchmarks that reflect real-world floating-point practice via a case study (Section˜6) that extracts 59 self-contained C floating-point benchmarks from functions in our dataset.
@@ 2. Related Work
Most works that focus on numerical programs are evaluated using different sets of hand-picked programs. Hand-picking programs can be good for stressing certain types of problems, but can just as easily—intentionally or not—suppress other problems and challenges. It is also unclear how to generalise from hand-picked problems to code “in the real world”. (Indeed, this was part of what prompted this research in the first place.)
@@
We are aware of only a single benchmark suite created specifically for floating-point analysis: FPBench (Damouche et al., 2017). FPBench is typically used to evaluate static analysis tools (Titolo et al., 2025). The FPBench benchmarks were collected from individual papers (Darulova et al., 2018; Solovyev et al., 2018) where they were hand-picked and are typically isolated arithmetic expressions. They are therefore not likely representative of challenges found in real-world programs. Techniques that target programs outside of this set choose their own set of benchmarks, e.g., programs that are larger (Das et al., 2020) or that use specific data structures (Isychev and Darulova, 2023).
@@
Some techniques are evaluated on a small number of hand-picked functions from numerical libraries, e.g., from the GNU Scientific Library (GSL) (Miao et al., 2024; Guo and Rubio-González, 2020; Zou et al., 2019, 2022), occasionally extended by some additional functions. For example, Di Franco et al. (Di Franco et al., 2017) study bug characteristics in numerical libraries (NumPy, LAPACK, GSL, etc.), and Liew et al. (Liew et al., 2017) have two teams independently hand-pick benchmarks for evaluating symbolic execution. In the HPC community, there are examples of picking a specific proxy application, such as Lulesh, as a representative of “HPC code” (Wang and Rubio-González, 2024a). The tool NSan (Courbet, 2021) is evaluated on the commercial SPECfp2006 benchmark set that contains scientific computing applications and that was designed for performance benchmarking. In what way these selected benchmarks are representative of general real-world code, e.g., beyond just library code or a specific application domain, is unclear and not discussed in these papers.
Real-World Usage through Code Studies
@@
We are not aware of any large-scale code study that investigates characteristics of real-world floating-point code. Previous works have investigated how different features are used in practice in different programming languages, for example in the context of static analysis for the R programming language (sihlerAnatomyRealWorldCode2024), dynamic features in Smalltalk (callauHowDevelopersUseDynamic2011), the use of eval and dynamic behaviour in JavaScript (richardsTheEvalThatMenDo2011; richardsAnAnalysisOfDynamicBehaviour2010), dynamic features or code complexity in Python (akerblomTracingDynamicFeatures2014; grotovComparisonPythonJupyter2022), inheritance, method chaining, and streams in Java (temperoHowDoJavaProgramsUseInheritance2008; dyerMiningBillionsAST2014; khatchadourian2020; keshkMethodChainingRedux2023), or gradual types in TypeScript and Python (troppmannTypedConfused2024). While this paper also studies features of code, these are sufficiently different that we cannot straightforwardly re-use existing infrastructure. We furthermore consider floating-point usage across many different programming languages.
While some features in dynamically typed programming languages such as Python can be detected statically (pengEmpiricalStudyCommon2021), others—specifically types—require programs to be run on suitable inputs (akerblomTracingDynamicFeatures2014). For example, to determine whether an addition (+) operates over floating-point values requires detection of the operand types at runtime. This cannot be straightforwardly automated; to run code it is necessary to obtain—for each project individually—all its required dependencies, and understand how to run the program in a representative way (including representative and meaningful input data and GUI interactions). Such information is not available in a consistent format that can be automatically processed at a large scale. We thus focus on statically typed programming languages only in this study.
GitHub Mining
@@
Our methodology draws inspiration from state-of-the-art practices in GitHub mining. Prior studies show that the majority of GitHub repositories are inactive, personal, or not software projects at all, and that most of the files are duplicates of others (kalliamvakouPromisesPerilsMining2014; lopesDejaVuMapCode2017; jupyterNotebooksonGitHub). Other work demonstrates that filtering repositories using extrinsic characteristics such as the number of stars introduces systematic biases in software repository studies (borgesWhatsGitHubStar2018; majFaultOurStars2024). In a reproduction study, Berger et al. reveal the risks of “careless” keyword searches (e.g., tagging commits as bug fixes if they contain the word "fix" without excluding "infix", "suffixes", or "does not fix" etc.) and emphasise the necessity of manual validation to ensure methodological soundness (berger_impact_2019).
@@
Several frameworks exist to simplify code studies on GitHub, but none fully satisfy our requirements. Queryable datasets mined from GitHub exist, yet they are either outdated (dyerBoaUltraLargeScale2015), affected by inconsistencies (gousios2012ghtorrent), or rely on GitHub stars (dabicSamplingProjectsGitHub2021; dabicSeartDataHub2024) that have been shown to bias results (majFaultOurStars2024). World of Code (maWorldCodeEnabling2021) continuously aggregates repositories from multiple sources and could be used to implement parts of our methodology, but it does not expose part of the metadata we require and does not allow downloading of full projects. Other frameworks target different aspects of repository analysis. CodeDJ (majCodeDJReproducibleQueries2021) ensures reproducible queries over projects by downloading full project histories, which exceeds our needs and would lead to an excessive amount of data being downloaded; the analysis would also still need to be implemented “from scratch”. GitCProc (casalnuovoGitcprocToolProcessing2017) enables keyword searches in commit histories but supports only a limited set of languages and does not exclude comments or string literals, which compromise validity. CodeQL (githubcodeql_2025) offers semantic analysis across multiple languages but requires full-project access, raising scalability concerns.
@@ 3. Methodology
Our goal is to study the prevalence and type of code operating on floating-point numbers in statically typed “real-world” code. To that end, we design a methodology that mainly follows the approach suggested by Maj et al. (majFaultOurStars2024) and that is designed in accordance with the ACM SIGSOFT Empirical Standards for repository mining (ralphEmpiricalStandardsSoftware2021). We first give an overview here, before discussing details in the next subsections.
3.1. Overview
We first define our target population: functions from non-trivial projects hosted on GitHub containing floating-point computations within code written in statically typed languages. GitHub alone contains an enormous amount of public repositories, and we assume here that they form a representative sample of real world software [^4] Although our methodology and implementation use GitHub as the primary data source, they are not specific to it. Both can be extended or adapted to other hosting platforms.. We restrict our study to statically typed code so that we can automate the analysis and thus to scale to significantly more projects.
Our final dataset consists of functions that use floating-point arithmetic. We focus on individual functions as most floating-point reasoning tools work on a per-function basis, though the corresponding files and projects are easily identified. Our quantitative analysis considers per-project, per-file and per-function properties, and so code not appearing in functions will be considered in the first two.
We cannot query GitHub for non-trivial projects containing such functions directly in part because each GitHub API call provides only limited information. We thus design a multi-step selection process or pipeline that progressively reduces the amount of data, increasing the level of granularity from the repository to the file and ultimately to the function level. An overview of the data selection pipeline is shown in Figure˜1.
@@ 3.1. Overview
Repository mining pipeline: sampling, metadata filtering, language selection, downloading, deduplication, and function extraction and analysis.
We begin by uniformly sampling public repositories on GitHub using GitHub’s Rest API (step 1), and by excluding forks and ‘trivial’ projects (defined in Section˜3.2, step 2 and 3). For each remaining repository, we query the list of used programming languages and retrieve the hash of the last commit on the main branch (step 4). We discard projects that do not have any files written in statically typed languages (step 5). At the end of this step, we have metadata about the repositories, after preliminary filtering.
Next, we download the content of the selected repositories and inspect individual files within these that are written in statically typed programming languages (step 6). For these we identify floating-point usage using keywords in the source code directly, e.g., focusing on type annotations (e.g., double and real) or mathematical functions (e.g., sin and exp). We discard data that we do not consider: files that are not written in statically typed languages, files that do not contain any keyword related to numerical computations, and files that are token-level duplicates of other files (in any considered repository, step 7). We then parse the remaining files, remove comments, and extract individual functions (step 8). We retain only functions that include keywords related to floating-point computations (step 9). In the final step we quantify usage of programming language constructs such as loops or conditional statements fully automatically based on parse trees.
The pipeline is designed to be fully automated, reproducible, modular and extensible. The sampling and filtering steps earlier in the pipeline are not specific to floating-point arithmetic and can be re-used for other code studies. We envision the floating-point specific dataset collected at the end of the pipeline to be used, e.g., to sample or validate benchmarks with specific features that are used for evaluating new floating-point automated reasoning techniques.